0.3kg of air at pressure of 350kN/m2 and a temperature of 35 oC receives heat energy at constant volume until its pressure become 700kN/m2. it then receives heat energy at constant pressure until its volume become 0.2289m3. determine the change of entropy during each process. take Cp = 1.006kJ/kgK and Cv = 0.717kJ/kgK
1
Expert's answer
2020-02-11T09:51:28-0500

"\delta Q=\Delta U+pdV=\frac{m}{M}C_VdT+pdV"




If "V=const"


"\delta Q=\Delta U+pdV=\frac{m}{M}C_VdT"


"\Delta S_1 =\int\frac{\delta Q}{T}=\int \frac{(m/M)C_VdT }{T}=\frac{m}{M}C_V\ln \frac{T_2}{T_1}=\frac{m}{M}C_V\ln \frac{p_2}{p_1}=\frac{0.3}{28.96\cdot 10^{-3}}\cdot 717\cdot\ln \frac{700\cdot 10^{-3}}{350\cdot 10^{-3}}=5148 J/K"




if "p=const"


"\delta Q=\Delta U+pdV=\frac{m}{M}C_VdT+pdV=\frac{m}{M}C_pdT"


"\Delta S_2 =\int\frac{\delta Q}{T}=\int \frac{(m/M)C_pdT}{T} = \frac{m}{M}C_p\ln \frac{T_2}{T_1}=\frac{m}{M}C_p\ln \frac{V_2}{V_1}"


"V_1=\frac{(m/M)RT_1}{p_1}=\frac{(0.3/0.02896)\cdot 8.31\cdot 308}{350\cdot 10^{3}}=0.07575m^3"


"\Delta S_2 =\frac{m}{M}C_p\ln \frac{V_2}{V_1}=\frac{0.3}{0.02896}
\cdot 1006\cdot \ln \frac{0.2289}{0.07575}=11524 J/K"


Answer. "\Delta S_1=5148 J/K; \Delta S_2=11524 J/K"